# Can Semantic Methods Enhance Team Sports Tactics? A Methodology for Football with Broader Applications

### v1

@@ Abstract
This paper explores how semantic-space reasoning, traditionally used in computational linguistics, can be extended to tactical decision-making in team sports. Building on the analogy between texts and teams — where players act as words and collective play conveys meaning—the proposed methodology models tactical configurations as compositional semantic structures. Each player is represented as a multidimensional vector integrating technical, physical, and psychological attributes; team profiles are aggregated through contextual weighting into a higher-level semantic representation.
Within this shared vector space, tactical templates such as high press, counterattack, or possession build-up are encoded analogously to linguistic concepts. Their alignment with team profiles is evaluated using vector-distance metrics, enabling the computation of tactical “fit” and opponent-exploitation potential. A Python-based prototype demonstrates how these methods can generate interpretable, dynamically adaptive strategy recommendations, accompanied by fine-grained diagnostic insights at the attribute level.
Beyond football, the approach offers a generalizable framework for collective decision-making and performance optimization in team-based domains—ranging from basketball and hockey to cooperative robotics and human–AI coordination systems. The paper concludes by outlining future directions toward real-world data integration, predictive simulation, and hybrid human–machine tactical intelligence.
@@ 1 Introduction
Despite significant progress in match analysis, a structural discontinuity persists between quantitative analytical tools—such as numerical performance indicators, spatial distributions, or xG-based predictive models—and qualitative contextual factors that critically influence collective performance. These include group cohesion, psychological resilience, team morale, residual energy, and the quality of tactical leadership [28, 15]. Current decision support systems (DSS) tend to emphasize easily measurable variables (e.g., physical metrics) while neglecting intangible dimensions that often prove decisive in dynamic, high-pressure contexts. This gap results in: (i) loss of strategically relevant information; (ii) limited adaptability and personalization of recommendations; and (iii) persistent reliance on subjective intuition in crucial phases of play [21].
To address these challenges, the present study introduces a Decision Support System for Tactical Optimization grounded in an innovative semantic-distance methodology. The key idea is to represent, within a common vector space, both the team’s contextual configuration — aggregating technical, physical, psychological, and organizational attributes — and the ideal profiles of canonical tactical strategies (e.g., high pressing, counterattack, possession build-up). The system then recommends the most coherent strategy by minimizing the semantic distance (computed via cosine and Euclidean metrics) between these two sets of vectors. This approach enables transparent integration of quantitative data and expert knowledge, allowing dynamic adaptation to evolving match contexts (e.g., changes in collective energy or morale).
A distinctive feature of this research lies in the transfer of a general semantic methodology— conceived initially to bridge analytical frameworks and decision heuristics—to the particular tactical domain of football. In the reference paper Recommending Actionable Strategies [5], semantic distance was used to connect abstract structures (such as the 6C model) with historical heuristic systems (like the Thirty-Six Stratagems), demonstrating the potential to mediate between conceptual traditions.
The present work extends this paradigm to the operational level of tactical practice, replacing:
  • general decision categories with 14 concrete macro-attributes capturing technical, psychological, and organizational dimensions of a team; and
  • general heuristics with a structured repertoire of canonical football strategies (e.g., high pressing, counterattack, positional defense).
Thus, the contribution goes beyond replication, applying semantic modeling to the tactical decision-making process to produce contextual, interpretable, and immediately actionable recommendations for coaching staff. This transposition—from general semantic theory to applied sports intelligence—represents the main innovative contribution of the work, aligning with the interdisciplinary research frontier that integrates NLP, decision theory, and sports science.
The main contributions of this research can be summarized as follows:
@@
    Formalization of an Integrated Semantic Model for Football – Adaptation and extension of the Recommending Actionable Strategies methodology to the football domain, defining 14 multidimensional macro-attributes that synthesize a team’s complexity (including Offensive Strength, Psychological Resilience, and Tactical Cohesion).
@@
    Development of a Tactical Recommender Prototype – Implementation, in Python, of a recommendation engine capable of aggregating heterogeneous data and comparing them against predefined tactical models. The prototype includes dynamic weighting mechanisms to adjust recommendations in real time (e.g., penalizing energy-intensive strategies when residual stamina is low).
@@
    Systematic Evaluation and Robustness Testing – Validation through both simulated scenarios (Advantage/Draw/Disadvantage $\times$ High/Low Energy) and retrospective analysis of real match data from German youth football. The evaluation includes ablation studies and robustness-to-noise analyses, ensuring the model’s interpretability and consistency across synthetic and empirical conditions.
In summary, this work contributes to the ongoing digital transformation of football by proposing an interpretable, flexible, and data-driven approach to tactical decision support that bridges the gap between numerical analytics and expert knowledge and lays the foundations for next-generation adaptive strategy systems.
The remainder of this paper is organized as follows. Section 2 reviews related work on performance analysis, decision support systems in sports, and semantic similarity methods. Section 3 presents the proposed methodology, including the formalization of the semantic model, the definition of macro-attributes, and the recommendation mechanism. Section 4 describes the prototype implementation in Python. Section 5 reports the experimental evaluation based on simulated scenarios, including ablation and robustness analyses. Section 6 extends the validation to real match data through retrospective case studies. Section 7 discusses the results, limitations, and practical implications. Finally, Section 8 concludes the paper and outlines directions for future work.
@@ 2.2 Canonical Tactical Strategies in Modern Football
Gegenpressing (Pressing After Loss). An aggressive evolution of pressing, aiming to recover the ball within 3–5 seconds after losing it by exploiting the opponent’s temporary disorganization. Extremely demanding, it requires maximal energy, readiness, and synchronization.
@@ 2.3 Semantic Distance Models
"With the advent of Transformer architectures (e.g., BERT, RoBERTa, Sentence-BERT) [3, 20], Contextual embeddings have dramatically improved representation quality, dynamically capturing meaning and outperforming static models such as Word2Vec and GloVe. These techniques have been widely adopted in information retrieval, question answering, text classification, and recommender systems [25].
@@ 3.2 Context Tree and Aggregation
    Root level: The 14 macro-attributes ($A_{1},\ldots,A_{14}$) that define the shared semantic space, computed by weighted combination of intermediate-level signals.
@@
All intermediate and final values are normalized to $[0,1]$ via min-max scaling against league or historical benchmarks, ensuring cross-team comparability.
Data Sources.
The context tree is designed to integrate multiple data streams:
  • Event data (e.g., Opta, StatsBomb): passes, shots, tackles, interceptions $\rightarrow$ technical attributes ($A_{1}$$A_{6}$, $A_{11}$, $A_{12}$)
  • Tracking data (e.g., SkillCorner, Second Spectrum): positions, velocities, distances $\rightarrow$ physical attributes ($A_{4}$, $A_{8}$, $A_{13}$)
  • Physiological monitoring (e.g., Catapult, Polar): heart rate, workload $\rightarrow$ energy and fatigue ($A_{8}$)
  • Qualitative assessments: coach ratings, historical stability $\rightarrow$ psychological/organizational attributes ($A_{7}$, $A_{9}$, $A_{14}$)
@@ 3.3 A Shared Semantic Space: 14 Macro-Attributes
Attribute Categories and Design Rationale.
The set of 14 attributes was designed to balance the following three design criteria, with Table 1 providing their complete specification:
  • Physical ($A_{8}$, $A_{12}$, $A_{13}$): Athletic resources and current energy state—what the team can sustain.
  • Psychological/Organizational ($A_{7}$, $A_{9}$, $A_{10}$, $A_{11}$, $A_{14}$): Intangible factors affecting collective performance—how the team responds under pressure.
Table 1: Complete specification of the 14 macro-attributes defining the shared semantic space.
@@
Technical/Tactical Dimensions
@@
Physical Dimensions
@@
Aggregation Functions.
@@
Dynamic vs. Static Attributes.
@@
  • Dynamic: $A_{8}$ (Residual Energy), $A_{9}$ (Team Morale), $A_{7}$ (Psychological Resilience under match stress)
  • Static: $A_{12}$ (Technical Base), $A_{13}$ (Physical Base), $A_{14}$ (Relational Cohesion)
@@
  • Context-dependent: $A_{5}$ (High Press Capability depends on energy), $A_{10}$ (Time Management becomes critical late in matches)
@@
This distinction informs the dynamic reweighting mechanism (Section 3.5), which adjusts attribute salience in response to evolving match conditions.

### v2

@@ Abstract
This paper explores how semantic-space reasoning, traditionally used in computational linguistics, can be extended to tactical decision-making in team sports. Building on the analogy between texts and teams—where players act as words and collective play conveys meaning—the proposed methodology models tactical configurations as compositional semantic structures. Each player is represented as a multidimensional vector integrating technical, physical, and psychological attributes; team profiles are aggregated through contextual weighting into a higher-level semantic representation. Within this shared vector space, tactical templates such as high press, counterattack, or possession build-up are encoded analogously to linguistic concepts. Their alignment with team profiles is evaluated using vector-distance metrics, enabling the computation of tactical “fit” and opponent-exploitation potential. A Python-based prototype demonstrates how these methods can generate interpretable, dynamically adaptive strategy recommendations, accompanied by fine-grained diagnostic insights at the attribute level. Evaluation through synthetic scenarios and a pilot study with real match data establishes internal consistency and feasibility of the approach; operational validity in live coaching contexts remains an open question for future prospective validation. Beyond football, the framework offers a potentially generalizable approach for collective decision-making in team-based domains—ranging from basketball and hockey to cooperative robotics and human–AI coordination systems. The paper concludes by outlining future directions toward real-world data integration, predictive simulation, and the validation work required before operational deployment
arXiv Note: This version corresponds to the final published article in Sci 2026, Vol. 8, Issue 3, Article 63. DOI: 10.3390/sci8030063. Formatted for arXiv compliance by removing publisher-specific formatting while preserving all scholarly content, corrections, figures, and references.
@@ 1 Introduction
Despite significant progress in match analysis, a fundamental disconnect persists between quantitative tools (performance indicators, spatial distributions, xG models) and qualitative factors that critically influence performance: group cohesion, psychological resilience, team morale, and residual energy [4, 5]. Current decision support systems emphasize easily measurable variables while neglecting intangible dimensions that often prove decisive under pressure.
Research gap: This disconnect causes (i) loss of strategically relevant information, (ii) limited adaptability of recommendations, and (iii) persistent reliance on subjective intuition in crucial match phases [6]. No existing framework provides a unified representation integrating quantitative performance data with qualitative contextual factors into a single, computationally tractable space for automated tactical recommendation.
To address this gap, the present study introduces a Decision Support System for Tactical Optimization using a novel semantic-distance methodology. The core innovation is representing both team states and tactical strategies as vectors in a shared 14-dimensional attribute space, enabling direct geometric comparison. The system recommends strategies by minimizing weighted Euclidean distance between a team’s profile and each strategy’s requirements; the rationale for this metric choice is detailed in Section 3.5.
This approach extends the semantic-distance methodology from Recommending Actionable Strategies [7]—originally developed for bridging analytical frameworks with decision heuristics—to operational football tactics. The key adaptations are:
  • replacing general decision categories with 14 concrete macro-attributes capturing technical, psychological, and organizational dimensions; and
  • replacing general heuristics with 20 canonical football strategies (e.g., high pressing, counterattack, positional defense).
What distinguishes this work: (i) unified representation integrating quantitative and qualitative factors; (ii) dynamic attribute weights adjusting in real time based on match context; (iii) transparent diagnostics showing why a strategy is recommended.
1.1 Objectives and Contributions
This study aims to: (1) formalize a semantic model encoding team states and strategies in a shared attribute space; (2) develop a prototype DSS with context-aware recommendations; (3) evaluate internal consistency and real-data feasibility.
The main contributions are:
@@ 1.1 Objectives and Contributions
    Semantic Model: 14 macro-attributes synthesizing team complexity, with 20 canonical strategies as ideal-profile vectors.
@@
    Adaptive Engine: Python prototype with dynamic weighting adjusting recommendations based on energy, time pressure, and opponent characteristics.
@@
    Systematic Validation: Evaluation through synthetic scenarios and German youth football data, including ablation and robustness analyses.
1.2 Paper Organization
Section 2 reviews related work; Section 3 presents methodology; Section 4 describes implementation; Section 5 reports evaluation; Section 6 presents the pilot study; Section 7 discusses limitations; Section 8 concludes.
@@ 2.2 Canonical Tactical Strategies in Modern Football
Gegenpressing (Pressing After Loss). An aggressive evolution of pressing, aiming to recover the ball within 3–5 s after losing it by exploiting the opponent’s temporary disorganization. Extremely demanding, it requires maximal energy, readiness, and synchronization.
@@ 2.3 Semantic Distance Models
“With the advent of Transformer architectures (e.g., BERT, RoBERTa, Sentence-BERT) [19, 20], Contextual embeddings have dramatically improved representation quality, dynamically capturing meaning and outperforming static models such as Word2Vec and GloVe. These techniques have been widely adopted in information retrieval, question answering, text classification, and recommender systems [21].
@@ 3.2 Context Tree and Aggregation
    Root level: The 14 macro-attributes ($A_{1},\ldots,A_{14}$) that define the shared semantic space, computed by a weighted combination of intermediate-level signals.
@@ 3.2.1 Aggregation Example
All intermediate and final values are normalized to $[0,1]$ via min-max scaling against league or historical benchmarks, ensuring cross-team comparability. The normalization procedure is specified in detail below.
3.2.2 Normalization Procedure
To ensure consistent scaling across teams and time periods, each raw attribute value $x$ is transformed to a normalized value $\tilde{x}\in[0,1]$ using the following protocol:
Min-Max Scaling Formula.
$\tilde{x}=\frac{x-x_{\min}}{x_{\max}-x_{\min}}$ (1)
where $x_{\min}$ and $x_{\max}$ are benchmark bounds derived from reference populations as specified below.
Benchmark Population Definition. Normalization benchmarks are computed from a reference population defined as follows:
  • League-level benchmarks (default): For professional deployments, benchmarks are derived from all players in the same league and division (e.g., Bundesliga, Serie A) over the reference window. This ensures that a normalized value of 0.5 represents league-average performance.
  • Competition-level benchmarks: For tournament contexts (e.g., Champions League, World Cup), benchmarks may be computed across all participating teams to reflect the elevated baseline.
  • Historical team benchmarks: For longitudinal tracking of a single team, benchmarks may be derived from that team’s own historical range, enabling detection of relative improvement or decline.
The current prototype uses synthetic benchmarks derived from the role-specific distributions in Table 22 (Appendix A), with $x_{\min}=\mu-2\sigma$ and $x_{\max}=\mu+2\sigma$ for each attribute–role combination.
Reference Window (Temporal Scope). To prevent instability from outlier matches, benchmarks are computed over a rolling window:
  • Season window (default): The most recent complete season (e.g., 34–38 matches for major European leagues). This captures stable population characteristics while remaining current.
  • Rolling window: For mid-season deployment, a rolling window of the most recent $N=10$ league matches provides more responsive benchmarks, updated weekly.
  • Fixed historical window: For retrospective analysis or cross-season comparison, a fixed reference period (e.g., 2022–23 season) ensures consistent scaling.
Clipping and Robustness. To handle outliers and ensure bounded outputs:
  1. 1.
    Pre-normalization clipping: Raw values outside $x_{\min},x_{\max}]$ are clipped to the boundary values before applying Equation (1). This prevents exceptional performances (positive or negative) from distorting the scale.
  2. 2.
    Robust benchmark estimation: Benchmarks may optionally use the 5th and 95th percentiles rather than true min/max to reduce sensitivity to extreme outliers:
    $x_{\min}=P_{5}(\mathcal{D}),\quad x_{\max}=P_{95}(\mathcal{D})$
    where $\mathcal{D}$ is the reference population distribution.
  3. 3.
    Floor for near-zero ranges: If $x_{\max}-x_{\min}<\epsilon$ (indicating near-constant values), the attribute is assigned the default value 0.5 to avoid division instability.
Leakage Prevention. In retrospective evaluation and real-time deployment, normalization must use only information available at decision time:
  1. 1.
    Temporal ordering: Benchmarks for match $t$ are computed from data up to match t1t-1 only. Future match data are never included in benchmark computation.
  2. 2.
    Held-out validation: When evaluating DSS performance over a test period, benchmarks are frozen at values computed from a prior training period. No benchmark updates occur during the test window.
  3. 3.
    Same-match exclusion: When computing benchmarks, the current match’s data are excluded to prevent self-referential scaling.
Current prototype scope: The implementation uses fixed synthetic benchmarks (role-specific $\mu\pm 2\sigma$) that do not require temporal updating, thereby avoiding leakage by construction. Production deployments should implement the rolling-window protocol with the temporal safeguards above.
3.2.3 Data Sources
The context tree is designed to integrate multiple data streams, each contributing to specific macro-attributes:
  • Event data (e.g., Opta, StatsBomb): passes, shots, tackles, interceptions $\rightarrow$ technical/tactical attributes ($A_{1}$$A_{6}$), tactical cohesion ($A_{11}$), technical base ($A_{12}$)
  • Tracking data (e.g., SkillCorner, Second Spectrum): positions, velocities, distances $\rightarrow$ transition speed ($A_{4}$), residual energy ($A_{8}$), physical base ($A_{13}$)
  • Physiological monitoring (e.g., Catapult, Polar): heart rate, workload $\rightarrow$ residual energy ($A_{8}$)
  • Qualitative assessments: coach ratings, historical stability $\rightarrow$ psychological attributes ($A_{7}$, $A_{9}$), organizational attributes ($A_{10}$, $A_{14}$)
@@ 3.2.3 Data Sources
3.2.4 Measurement Framework
Integrating heterogeneous data streams requires explicit policies for handling missing data, quality control, and uncertainty propagation. The DSS adopts the following minimal measurement framework:
Data Stream Reliability Tiers. Each data source is assigned a reliability tier reflecting measurement precision and update frequency:
  • Tier 1 (High reliability): Event data from professional providers (Opta, StatsBomb)—validated, near-complete, low latency. Assigned confidence weight $c_{1}=1.0$.
  • Tier 2 (Medium reliability): Tracking data (SkillCorner, Second Spectrum)—high precision but potential occlusion gaps; physiological monitoring (Catapult, Polar)—device-dependent accuracy. Assigned $c_{2}=0.85$.
  • Tier 3 (Lower reliability): Qualitative assessments (coach ratings, historical proxies)—subjective, infrequently updated. Assigned $c_{3}=0.70$.
Missing Data Policy. When an input variable is unavailable, the following hierarchy applies:
  1. 1.
    Imputation from correlated sources: If a higher-tier source for the same attribute exists, use it with adjusted confidence. For example, if physiological $A_{8}$ data are missing, estimate from tracking-derived distance covered.
  2. 2.
    Historical baseline: If no current-match data exist, use the team’s season average for that attribute, flagged with reduced confidence ($c\leftarrow 0.5\cdot c_{\text{tier}}$).
  3. 3.
    Neutral default: If no historical data exist, assign the attribute the midpoint value (0.5) with minimal confidence ($c=0.3$), ensuring the attribute contributes little to distance until better data arrive.
Temporal Alignment. Data streams operate at different frequencies: event data are discrete (per-action), tracking data are high-frequency (25 Hz), and qualitative assessments are episodic (pre-match, halftime). The aggregation module aligns all inputs to a common temporal frame:
  • Match phase granularity: Attributes are computed per phase (first half, second half) or per 15-min window for finer resolution.
  • Windowed aggregation: High-frequency tracking data are averaged over the alignment window; event data are accumulated.
  • Carry-forward for episodic inputs: Qualitative assessments persist until updated (e.g., halftime morale rating carries into second half unless revised).
Quality Control. Basic outlier detection is applied before aggregation:
  • Player-level attributes outside $\mu-3\sigma,\mu+3\sigma]$ (based on role-specific distributions) are flagged and clamped to boundary values.
  • Implausible physiological readings (e.g., heart rate ¡ 40 or ¿220) are discarded and imputed from recent history.
  • Event data with missing location or timestamp fields are excluded from spatial aggregations but retained for count-based metrics.
Uncertainty Propagation and Reporting. Rather than reporting point-estimate distances, the DSS can optionally compute distance intervals reflecting input uncertainty:
$d_{\text{interval}}=\left[d(V_{\text{team}}^{-},V_{\text{strategy}}),\;d(V_{\text{team}}^{+},V_{\text{strategy}})\right]$ (2)
where $V_{\text{team}}^{-}$ and $V_{\text{team}}^{+}$ are lower and upper bounds on the team vector, derived by perturbing each attribute $A_{j}$ by $\pm\epsilon_{j}$ proportional to $(1-c_{j})$, where $c_{j}$ is the confidence weight for that attribute’s data source. When intervals for competing strategies overlap, the DSS flags the recommendation as uncertain and presents the top-$k$ alternatives rather than a single choice.
Recommendation Stability Under Measurement Error. The robustness analysis in Section 5 (Monte Carlo perturbations, $\pm 5\%$ noise, $N=100$) provides an empirical estimate of recommendation stability: 89.3% top-1 consistency across plausible measurement error. This serves as the baseline stability metric; deployments with lower-tier data should expect reduced consistency and should increase the perturbation range accordingly (e.g., $\pm 10\%$ for Tier 3–dominant profiles).
Current prototype scope: The implementation assumes Tier 1 data availability (complete event data) with synthetic generation for missing streams. The full measurement framework described above is designed for production deployment; the robustness analyses in Section 5 simulate its behavior under controlled noise conditions.
@@ 3.3 A Shared Semantic Space: 14 Macro-Attributes
3.3.1 Complete Attribute Set
The semantic space comprises exactly 14 macro-attributes, indexed $A_{1}$ through $A_{14}$ with no gaps. Table 1 presents the complete ordered list; Table 2 provides detailed specifications grouped by functional category.
Table 1: Sequential enumeration of all 14 macro-attributes ($A_{1}$$A_{14}$).
ID Attribute Name Category Variability
$A_{1}$ Offensive Strength Technical/Tactical Static
$A_{2}$ Defensive Strength Technical/Tactical Static
$A_{3}$ Midfield Control Technical/Tactical Static
$A_{4}$ Transition Speed Technical/Tactical Semi-dynamic
$A_{5}$ High Press Capability Technical/Tactical Context-dependent
$A_{6}$ Width Utilization Technical/Tactical Static
$A_{7}$ Psychological Resilience Psychological Dynamic
$A_{8}$ Residual Energy Physical Dynamic
$A_{9}$ Team Morale Psychological Dynamic
$A_{10}$ Time Management Organizational Context-dependent
$A_{11}$ Tactical Cohesion Organizational Semi-dynamic
$A_{12}$ Technical Base Physical Static
$A_{13}$ Physical Base Physical Static
$A_{14}$ Relational Cohesion Organizational Static
Table 2: Detailed specification of macro-attributes with definitions and aggregation sources.
@@ 3.3.1 Complete Attribute Set
Technical/Tactical Dimensions ($A_{1}$$A_{6}$)
@@
Psychological/Physical Dimensions ($A_{7}$$A_{9}$, $A_{12}$$A_{13}$)
Table 3: Cont.
ID Attribute Name Definition & Aggregation Source
@@
3.3.2 Design Rationale
The 14-attribute set was designed according to the following principles:
  1. 1.
    Completeness: The set covers the major dimensions of team performance identified in sports science literature: technical skill, tactical capability, physical capacity, psychological state, and organizational coherence.
  2. 2.
    Orthogonality: Attributes were selected to minimize redundancy. For example, $A_{1}$ (Offensive Strength) captures goal-scoring capability, while $A_{12}$ (Technical Base) captures underlying skill level—a team may have high technical quality but poor offensive output due to tactical misalignment.
  3. 3.
    Measurability: Each attribute can be estimated from available data sources: event data for $A_{1}$$A_{6}$, $A_{11}$, $A_{12}$; tracking data for $A_{4}$, $A_{8}$, $A_{13}$; physiological monitoring for $A_{8}$; and qualitative assessment for $A_{7}$, $A_{9}$, $A_{14}$.
  4. 4.
    Tactical relevance: Each attribute has clear implications for strategy selection. For instance, low $A_{8}$ (energy) constrains high-pressing options; high $A_{4}$ (transition speed) enables counterattacking; strong $A_{11}$ (cohesion) supports complex positional play.
3.3.3 Attribute Categories
For exposition, we group attributes into three functional categories (though the DSS treats all 14 dimensions uniformly in distance computations):
  • Psychological/Physical ($A_{7}$$A_{9}$, $A_{12}$$A_{13}$): Individual and collective resources—what the team can sustain.
  • Organizational ($A_{10}$, $A_{11}$, $A_{14}$): Coordination and adaptation capabilities—how the team functions as a unit.
Table 2 provides the complete specification with aggregation sources for each attribute.
3.3.4 Aggregation Functions
@@
3.3.5 Dynamic vs. Static Attributes
@@
  • Dynamic: $A_{7}$ (Psychological Resilience), $A_{8}$ (Residual Energy), $A_{9}$ (Team Morale)—vary significantly during a match based on events and fatigue.
  • Static: $A_{1}$$A_{3}$, $A_{6}$, $A_{12}$$A_{14}$—determined by squad composition; stable within a match.
@@
  • Context-dependent: $A_{4}$ (Transition Speed), $A_{5}$ (High Press Capability), $A_{10}$ (Time Management), $A_{11}$ (Tactical Cohesion)—baseline is static but effective value depends on match context (e.g., $A_{5}$ is constrained by $A_{8}$; $A_{10}$ becomes critical late in matches).
@@
The variability classification for all 14 attributes is summarized in Table 1. This distinction informs the dynamic reweighting mechanism (Section 3.5), which adjusts attribute salience in response to evolving match conditions.
3.3.6 Construct Validity: Input Overlap and Multicollinearity
Several macro-attributes share underlying player-level inputs, creating potential redundancy. Specifically:
  • $A_{7}$ (Psychological Resilience) and $A_{9}$ (Team Morale) both aggregate resilience and aggression with similar weights (0.7/0.3 vs. 0.6/0.4).
  • $A_{8}$ (Residual Energy) uses stamina and resilience, sharing the latter with $A_{7}$ and $A_{9}$.
  • $A_{3}$ (Midfield Control) and $A_{6}$ (Width Utilization) both aggregate xA from fullbacks and central midfielders.
To quantify the resulting correlations, we generated 500 synthetic teams using the player attribute distributions in Table 22 (Appendix A) and computed the pairwise correlation matrix. Table 4 reports correlations for the most affected attributes.
Table 4: Pairwise correlations among attributes with overlapping inputs ($n=500$ synthetic teams).
$A_{3}$ $A_{6}$ $A_{7}$ $A_{8}$ $A_{9}$
$A_{3}$ (Midfield Control) 1.00 0.90 0.05 0.00 0.05
$A_{6}$ (Width Utilization) 0.90 1.00 0.04 0.15 0.04
$A_{7}$ (Psych. Resilience) 0.05 0.04 1.00 0.32 0.98
$A_{8}$ (Residual Energy) 0.00 0.15 0.32 1.00 0.26
$A_{9}$ (Team Morale) 0.05 0.04 0.98 0.26 1.00
Two attribute pairs exhibit high correlations: $A_{7}$$A_{9}$ ($r=0.98$) and $A_{3}$$A_{6}$ ($r=0.90$). The corresponding Variance Inflation Factors (VIF) are 35.1 ($A_{7}$), 37.0 ($A_{9}$), 21.6 ($A_{3}$), and 18.4 ($A_{6}$)—well above the conventional threshold of 10, indicating severe multicollinearity. The remaining 10 attributes have VIF <5<5, suggesting acceptable independence.
3.3.7 Implications for Distance Computation
Under Euclidean distance, correlated attributes contribute partially redundant information, effectively double-counting shared variance. For the $A_{7}$$A_{9}$ pair, a team’s psychological profile influences distance along two nearly parallel axes, inflating the contribution of resilience/aggression relative to other dimensions.
3.3.8 Design Justification
Despite the statistical correlation, we retain both $A_{7}$ and $A_{9}$ (and both $A_{3}$ and $A_{6}$) for the following reasons:
  1. 1.
    Conceptual distinctness: In sports psychology, resilience (ability to recover from setbacks) and morale (current motivational state) are treated as related but distinct constructs [25]. A team may have high baseline resilience yet low in-match morale due to recent conceded goals. Similarly, midfield control (tempo dictation) and width utilization (flank exploitation) represent tactically distinct capabilities that happen to draw on overlapping personnel.
  2. 2.
    Strategy vector differentiation: Tactical templates assign different weights to these correlated attributes. For example, “High Press” requires high $A_{7}$ but is neutral on $A_{9}$, while “Cautious Horizontal Play” prioritizes $A_{9}$ (maintaining composure) over $A_{7}$ (bouncing back from pressure). The correlation at the team level does not imply identical strategic relevance.
  3. 3.
    Dynamic divergence: Under match conditions, $A_{7}$ and $A_{9}$ can diverge: morale ($A_{9}$) is modulated by score state and momentum, while resilience ($A_{7}$) reflects a more stable trait. The current prototype does not fully exploit this divergence, but the architectural separation enables future refinement.
3.3.9 Mitigation Strategies
We acknowledge that the current Euclidean metric does not account for attribute covariance. Several mitigation approaches are available:
  • Mahalanobis distance: Replacing Euclidean with Mahalanobis distance, $d_{M}(x,y)=\sqrt{(x-y)^{\top}\Sigma^{-1}(x-y)}$, would down-weight correlated dimensions automatically. However, this requires estimating the covariance matrix $\Sigma$ from representative team data, which is unavailable for the current prototype.
  • Principal component projection: Projecting the 14-dimensional space onto principal components would decorrelate the axes. This sacrifices interpretability (components are linear combinations rather than named attributes) but may be appropriate for purely predictive applications.
  • Attribute consolidation: Merging $A_{7}$$A_{9}$ into a single “Psychological State” dimension and $A_{3}$$A_{6}$ into “Midfield Effectiveness” would reduce redundancy but lose the conceptual granularity valued by coaching staff.
  • Regularized weighting: Applying lower dynamic weights to correlated pairs (e.g., halving $w_{7}$ and $w_{9}$) would reduce their combined influence, approximating the effect of Mahalanobis correction.
For the current prototype, we retain the 14-attribute Euclidean formulation with the following justification: (i) the DSS is designed for interpretability, and named attributes are more actionable than principal components; (ii) the high-correlation pairs ($A_{7}$$A_{9}$, $A_{3}$$A_{6}$) represent 4 of 14 dimensions, limiting the overall inflation; and (iii) the dynamic weighting mechanism (Section 3.6.2) already modulates attribute salience, partially mitigating fixed-correlation effects. Future work should implement Mahalanobis distance once sufficient real-world team data are available to estimate $\Sigma$ reliably.
3.3.10 Empirical Impact Assessment
To quantify the practical effect of multicollinearity on strategy recommendations, we conducted a sensitivity analysis comparing Euclidean rankings with a correlation-adjusted baseline. Specifically:
  1. 1.
    We computed strategy rankings for 100 synthetic team profiles using standard Euclidean distance.
  2. 2.
    We repeated the analysis using a “consolidated” 12-attribute space where $A_{7}$$A_{9}$ and $A_{3}$$A_{6}$ were each merged into single dimensions (averaging their values).
  3. 3.
    We measured rank correlation (Kendall’s $\tau$) between the two ranking schemes.
The mean rank correlation was $\tau=0.91$ (95% CI: 0.87–0.94), indicating that the top-ranked strategies are largely stable despite the redundancy. The primary effect of consolidation was minor reordering among middle-ranked strategies with similar distances. Critically, the top-1 recommendation matched in 94% of cases, and the top-3 set matched in 89% of cases.
These results suggest that while the multicollinearity is statistically significant, its practical impact on DSS recommendations is limited. Nevertheless, the full 14-dimensional correlation matrix is provided in Appendix A.7 for transparency, and we recommend that future deployments with real team data implement Mahalanobis distance or apply the regularized weighting correction described above.